Solve the system by addition.
4x – 4y = 9
8x + 2y = –7

Answer by sdmmadam@yahoo.com(530)

Solve the system by addition.
4x – 4y = 9 ----(1)
8x + 2y = –7 ----(2) X2
(to equalise the coefficients of y (numerically) multiply (2) through out by 2
16x+4y = -14 ----(3)
4x – 4y = 9 ----(1)
(3) + (1) implies
(16x+4x) +(4y-4y)= (-14+9) (adding left with left and right with right)
20x +0 = -5
20x = -5
x = -5/20 = -1/4
Putting x=-1/4 in (1)
4x-4y = 9
4X(-1/4) - 4y = 9
-1-4y = 9
-1-9 =4y
-10 = 4y
y = (-10)/4 = -5/2
Answer: x =(-1/4) and y = (-5/2)
Verification: putting x =(-1/4) and y = (-5/2)in 8x + 2y = –7 ----(2)
LHS = 8x+2y = 8X(-1/4)+2X(-5/2) = -2-5 = -7 =RHS
Hence our values are right
Note:You may divide (1) through out by 2 and write it as 2x – 2y = 9/2----(*) and then add it(*) up with (2)which is 8x + 2y = –7 ----(2)giving
(2x+10x)= (9/2-7)
10x = -5/2 which givis x=-5/4 and therefore y = -5/2